Calculate the concentration of NH3 required dissolving 0.01 mol of AgCl in a liter of
solution. Ksp(AgCl) =1.56x10-10 Kdiss (Ag (NH3)2+) =6.8x10-8 ?
1
Expert's answer
2020-09-24T13:50:22-0400

C(AgCl) = 0.01 mol/1.0L = 0.01 M

AgCl(s) + 2NH3 <=> [Ag(NH3)2]+ + Cl-

0.01 M 2×0.01 M 0.01 M 0.01 M

AgCl <=> Ag+ + Cl-

[Ag+] = [Cl-]

Ag + 2NH3 (aq) → [Ag(NH3)2+] (aq)

"k_{dis}= \frac{[NH_3]^2[Ag^+] }{ [[Ag(NH_3)_2^+]]} = 6.8×10^{-8}"

Ksp = [Ag+][Cl-] = [Ag+]2 = 1.56×10-10

[Ag+] = 1.25×10-5

"6.8×10^{-8} = \frac{[NH_3]^2(1.25×10^{-5}) }{ 0.01}"

[NH3]2 = 5.44×10-5

[NH3] = 7.37×10-3 M

[NH3]total = 7.37×10-3 + (2×0.01) M = 0.0273 M